Find the area of a triangle whose altitude is 7 feet and whose base is 22 feet
Answer by bucky(2189)

The area of a triangle is found by using the formula:
.
A = (1/2) * b * h
.
where A is the area, b is the length of the base, and h is the height or the altitude of
the triangle.
.
We must stay in the same system of units. In this problem the base and altitude are in
feet so the area will be in square feet.
.
All you have to do now is substitute the values into the equation to get:
.
A = (1/2) * 22 * 7
.
Since (1/2) times 22 is 11, the area is:
.
A = 11 * 7 = 77 square feet
.
Hope this helps you to understand the problem and how to solve it.